An object falls from a high building and hits the ground in 9.0 seconds ignoring air assistance what is the distance that it fell


1
Expert's answer
2022-05-05T13:56:55-0400
"H=\frac{gt^2}{2}=\frac{9.8*9.0^2}{2}=397\:\rm m"